OpenStudy (anonymous):

A father is 1 year younger than 3 times the age of his son. The son is 20 years old. Let F = the age of the father. Which formula would calculate the father's age?

OpenStudy (anonymous):

(20 times 3)-1

OpenStudy (anonymous):

F = 3s - 1 if s = son's age F = 3(20) - 1 = 59.